A small town council must decrease operating costs by 17%.The present operating cost is $476,000. The council reduces the operating cost to $397,000. Is that enough? Explain.
Answer by Mathtut(3670)

476000-476000(.17)=476000-80920=395080
:
therefore, although close ,they did not reduce by quite enough.